This coming summer, we are meeting up with some other family members at Family Fiesta Amusement Park. The ticket price for children is $17 and for adults the ticket price is $31. The total price for all 20 of us to enter the park is $438. b. Write an equation, using A for adults and C for children, that illustrates how much money was spent.
Answer by rfer(16322)

a+c=20 a=20-c
31a+17c=438
31(20-c)+17c=438
620-31c+17c=438
-14c=-182
c=13 children
a=7 adults